OpenStudy (anonymous):

Rewrite with positive exponents. Assume that even roots are of nonnegative quantities and that all denominators are nonzero. (3y(2/3))^-3

OpenStudy (bahrom7893):

(3^(-3)y^(-2)) = 1/(27y^2)

OpenStudy (anonymous):

this is what i got so tell me did i do it wrong

OpenStudy (anonymous):

3^(-3) y^(-3 * 2/3) = 3^(-3) y^(-2) = .1 . . .1 . . . . . . . 1 ---- x ---- . = . ---------- 3^3 . y^2 . . . . 3(3y)^2

OpenStudy (bahrom7893):

that's correct btw

OpenStudy (anonymous):

ok thank u so much